OpenStudy (ravenxcx):

Algebra 2 Help? (Connections Academy) 1.) the graph of g(x) is f(x) translated to the left 7 units and down 3 units. What is the function rule for g(x) given f(x) = x^2? 2.) for g(x)= |3x+2| + 6, find g(-4).

OpenStudy (anonymous):

1.) g(x) = f(x+7)-3 Formula for horizontal translations: g(x)=f(x-h) h being the distance translated, If the translation is right-substitute h with the positive number; if the translation is left-substitute h with the negative number. Formula for vertical translations: g(x)= f(x)-k k being the distance translated, If the translation is up- substitute k with the positive number; if the translation is down-substitute k with the negative number... Hope this helps! 2.) g(-4)=|3x+2|+6 g(-4)=|3(-4)+2|+6 g(-4)=|-12+2|+6 g(-4)=|-10|+6 g(-4)=10+6 g(-4)=16 Just remember for g(x) the x in parenthesis is the x in the equation, so you can just substitute and then solve.